OpenStudy (anonymous):

Medal to correct answer! Solve the equation. 3(4-2x) = -2x

OpenStudy (anonymous):

please help me plz plz plz

HanAkoSolo (jamierox4ev3r):

alright… so first here you would distribute. DO you understand how to do that?

HanAkoSolo (jamierox4ev3r):

here, I'll just give you the process rlly quick… 3(4-2x)=-2x 12-6x=-2x +6x +6x -------- 12=4x x=3

HanAkoSolo (jamierox4ev3r):

let me know if you have any questions, but I have to go right now :D